# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

### v2

@@ Abstract
Determining the complexity of computing Gröbner bases is an important problem both in theory and in practice, and for that the solving degree plays a key role. In this paper, we study the solving degrees of affine semi-regular sequences and their homogenized sequences. Some of our results are considered to give mathematically rigorous proofs of the correctness of methods for computing Gröbner bases of the ideal generated by an affine semi-regular sequence. This paper is a sequel of the authors’ previous work [30] and gives additional results on the solving degrees and important behaviors of Gröbner basis computation.
We also define the generalized degree of regularity for a sequence of homogeneous polynomials. For the homogenization of an affine semi-regular sequence, we relate its generalized degree of regularity with its maximal Gröbner basis degree (i.e., the solving degree of the homogenized sequence). The definition of a generalized (cryptographic) semi-regular sequence is also given, and it derives a new cryptographic assumption to estimate the security of cryptosystems and signature schemes. From our experimental observation, we raise a conjecture and some questions related to this generalized semi-regularity. These new definitions and our results provide a theoretical formulation of (somehow heuristic) discussions done so far in the cryptographic community.
@@ 1 Introduction
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. We denote by $\mathbb{A}_{K}^{n}$ (resp. $\mathbb{P}_{K}^{n}$) the $n$-dimensional affine (resp. projective) space over $K$. Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over $K$. For a given monomial ordering $\prec$ on the set of monomials in $R$, let $\mathrm{LM}(f)$ denote the leading monomial of $f\in R\smallsetminus\{0\}$ with respect to it. For a non-empty subset $F\subset R\smallsetminus\{0\}$, put $\mathrm{LM}(F):=\{\mathrm{LM}(f):f\in F\}$. A set $F$ (resp. a sequence $\bm{F}$) of polynomials in $R$ is said to be homogeneous if the elements of $F$ (resp. $\bm{F}$) are all homogeneous, and otherwise $F$ is said to be affine. We denote by $\langle F\rangle_{R}$ (or $\langle F\rangle$ simply) the ideal generated by a non-empty subset $F$ of $R$. For a polynomial $f$ in $R\smallsetminus\{0\}$, let $f^{\rm top}$ denote its maximal total degree part which we call the top part of $f$, and let $f^{h}$ denote its homogenization in $R^{\prime}=R[y]$ by an extra variable $y$, see Subsection A.2 below for details. For a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus\{0\})^{m}$, we also set $\bm{F}^{\rm top}:=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ and $\bm{F}^{h}:=(f_{1}^{h},\ldots,f_{m}^{h})$. For a finitely generated graded $R$-(or $R^{\prime}$-)module $M$, we also denote by ${\rm HF}_{M}$ and ${\rm HS}_{M}$ its Hilbert function and its Hilbert–Poincaré series, respectively.
@@
A Gröbner basis of an ideal $I$ in $R$ is defined as a special kind of generating set for $I$, and it gives a computational tool to determine many properties of $I$. A typical application of computing Gröbner bases is solving the multivariate polynomial (MP) problem: Given a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of $m$ polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus\{0\}$, find $(a_{1},\ldots,a_{n})\in K^{n}$ such that $f_{i}(a_{1},\ldots,a_{n})=0$ for all $i$ with $1\leq i\leq m$. A particular case where $f_{1},\ldots,f_{m}$ are all quadratic is called the MQ problem, and its hardness is applied to constructing public-key cryptosystems and digital signature schemes that are expected to be quantum resistant. Therefore, analyzing the complexity of computing Gröbner bases is one of the most important problems both in theory and in practice.
@@
An algorithm for computing Gröbner bases was proposed first by Buchberger [6], and so far a number of its improvements such as the $F_{4}$ [18] and $F_{5}$ [19] algorithms have been proposed. In determining the complexity of computing Gröbner bases, as we will see in the first paragraph of Subsection 2.2 below, one of the most important cases is the case where the input system is zero-dimensional and where the monomial ordering is graded (i.e., degree-compatible), and we focus on that case in the rest of this paper. Namely, we suppose that the input sequence $\bm{F}=(f_{1},\ldots,f_{m})$ admits a finite number of zeros in $\mathbb{A}^{n}_{\overline{K}}$ (resp. $\mathbb{P}^{n-1}_{\overline{K}}$) if $\bm{F}$ is affine (resp. homogeneous), and we consider a monomial ordering $\prec$ on $R$ that compares monomials first by their total degrees, e.g., a degree reverse lexicographical (DRL) ordering. Then, the complexity of the Gröbner basis computation for $F=\{f_{1},\ldots,f_{m}\}$ is estimated as a function of the solving degree(s): To the authors’ best knowledge, there are three (in fact four) kinds of definitions of solving degree, and they will be rigorously described in Subsection 2.2 below. In the first definition, the solving degree is defined as the highest degree of the polynomials involved during the Gröbner basis computation. Since this solving degree depends on an algorithm $\mathcal{A}$ that one adopts, we denote it by $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$. On the other hand, in the second and the third definitions, which were originally provided in a series of Gorla et al.’s studies (cf. [8], [4], [24], [9], [23]), we can see that the solving degrees do not depend on an algorithm, but only on $F$ and $\prec$. The solving degree in the second (resp. third) definition is defined by using Macaulay matrices (resp. those with mutants), and it is denoted by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ (resp. $\mathrm{sd}_{\prec}^{\rm mut}(F)$) in this paper, where the subscripts “mac” and “mut” stand for Macaulay matrices and mutants respectively. Note that, when $F$ is homogeneous, these three solving degrees coincide with one another (for $\mathcal{A}$ with suitable setting) and we call them the solving degree simply; they are equal to the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec}(F)$ of $F$ with respect to $\prec$. In this case, we can apply a well-known bound [31, Theorem 2] by Lazard. In the following, we mainly treat with the case where $F$ is affine.
@@
  • ([8, Theorem 11]) When $K=\mathbb{F}_{q}$, the solving degree $\mathrm{sd}_{\prec}^{\rm mac}(F)$ for a DRL ordering $\prec$ can be upper-bounded by the Macaulay bound $d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell=\min\{n+1,\ell\}$, if $F$ contains the field equations $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
@@
  • Semaev-Tenti [40] (see also Tenti’s PhD thesis [41]) constructed a Buchberger-like algorithm $\mathcal{A}$ for the case $K=\mathbb{F}_{q}$ such that $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$ with $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$ for a DRL ordering $\prec$, assuming that $\{x_{i}^{q}-x_{i}:1\leq i\leq n\}\subset F$ and $\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D$. Here $d_{\rm reg}(\langle F^{\rm top}\rangle)$ is the degree of regularity for $F^{\rm top}$, i.e., the smallest integer $d$ with $R_{d}=\langle F^{\rm top}\rangle_{d}$, where $R_{d}$ denotes the homogeneous part (component) of degree $d$ and where we set $I_{d}=I\cap R_{d}$ for a homogeneous ideal $I$ of $R$.
  • Caminata-Gorla proved in [9, Theorem 3.1] that $\mathrm{sd}_{\prec}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\prec}}(F)\}$ for any graded monomial ordering $\prec$, where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]). Recently, Salizzoni [39] also proved $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq D+1$, in the case where $\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D<\infty$.
In this paper, by a mathematically rigorous way following Gorla et al.’s works, we study the solving degrees and related Gröbner bases of affine semi-regular polynomial sequences, where a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of (not necessarily homogeneous) polynomials is said to be affine semi-regular (resp. affine cryptographic semi-regular) if $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is semi-regular (resp. cryptographic semi-regular), see Definitions 2.1.3, 2.1.9, and 2.1.12 for details. Note that homogeneous semi-regular sequences are conjectured by Pardue [36, Conjecture B] to be generic sequences of polynomials (see e.g., [36] for the definition of genericness), and affine (cryptographic) semi-regular sequences are often appearing in the construction of multivariate public key cryptosystems and digital signature schemes. As a sequel of the authors’ previous work [30], we investigate further results on the solving degrees and on behaviors of the computation of Gröbner bases.
@@
Theorem 1 (Theorem 3.1.1, Remark 3.1.2, Remark 3.1.3 and Corollary 3.1.5).
@@
    There exists $d_{0}$ such that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for all $d$ with $d\geq d_{0}$, namely the number of projective zeros for $F^{h}$ is finite at most. Here, $d_{0}$ gives an upper-bound on the solving degree of $F^{h}$ (or equivalently the maximal degree of the Gröbner basis of $\langle F^{h}\rangle$).
@@
    ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\prod_{i=1}^{m}(1-z^{d_{i}})/(1-z)^{n+1}\pmod{z^{D}}$, so that $F^{h}$ is $D$-regular, equivalently $\mathrm{syz}(F^{h})_{<D}=\mathrm{tsyz}(F^{h})_{<D}$. Here we denote by $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$ the module of syzygies of $F^{h}$ and that of trivial syzygies of $F^{h}$, respectively (see Appendix A.1 for the definition of $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$).
As for (3) of Theorem 1, it follows from the proof of Lazard’s bound [31, Theorem 2] that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$ for a DRL ordering $\prec$ (we give an explicit proof for this in Lemma 2.2.2 below), where $\prec^{h}$ is the homogenization of $\prec$. As in [8, Theorem 11] recalled above, we can apply Lazard’s bound to obtaining $d_{0}\leq d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{i}=\deg(f_{i})$ and $\ell=\min\{n+1,\ell\}$, assuming $d_{1}\geq\cdots\geq d_{m}$ in descending order. As an additional result in this paper, we also obtain the following upper-bound on the solving degree of $F^{h}$:
@@
    Let $S_{0}$ be the saturation exponent of $(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$, that is, the minimum integer $s$ such that $(\langle F^{h}\rangle:\langle y^{s}\rangle)=(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$. Then the solving degree of $F^{h}$ is upper-bounded by $D+s_{0}$.
Based on Theorem 1, we can explore the computations of reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ in Section 4 below, dividing the cases into the degree less than $D$ or not. More precisely, denoting by $G$, $G_{\rm hom}$, and $G_{\rm top}$ the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$, we revisit [30, Section 5] and obtain more precise results:
Theorem 3 (Section 4; cf. [30, Section 5]).
@@
Furthermore, based on Lemma 2.2.2 below, for homogeneous ideals of projective dimension 0, we naturally extend the notion of degree of regularity. The new notion of generalized degree of regularity plays a very important role on the computation of Gröbner bases for such ideals, see Subsections 2.3 and 4.3. The following proposition summarizes theoretical results on this generalized degree of regularity:
Proposition 1 (Proposition 2.3.3).
@@
    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
@@
Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
Finally, in Subsection 4.3, we give some observation on the behavior of Gröbner basis computation based on our experiments, from which we arrive at a conjecture on the Hilbert–Poincaré series for affine polynomial sequences without constant terms. For this conjecture, we generalize the notion of cryptographic semi-regular in Subsection 2.3.
@@
  • ${\rm HF}_{M}$: The Hilbert function of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HF}_{M}(d)=\mathrm{dim}_{K}M_{d}$ for each $d\in{\color[rgb]{0,0,0}\mathbb{Z}_{\geq 0}}$.
  • ${\rm HS}_{M}$: The Hilbert–Poincaré series of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HS}_{M}(z)=\sum_{d=0}^{\infty}{\rm HF}_{M}(d)z^{d}\in\mathbb{Z}\llbracket z\rrbracket$.

### v3

@@ Abstract
Determining the complexity of computing Gröbner bases is an important problem both in theory and in practice, and for that the solving degree plays a key role. In this paper, we study the solving degrees for affine semi-regular sequences and their homogenized sequences. Some of our results are considered to give mathematically rigorous proofs of the correctness of methods for computing Gröbner bases of the ideal generated by an affine semi-regular sequence. This paper is a sequel of the authors’ previous work [31] and gives additional results on the solving degrees and important behaviors of Gröbner basis computation.
We also define the generalized degree of regularity for a sequence of homogeneous polynomials. For the ideal generated by the homogenization of an affine semi-regular sequence, we relate its generalized degree of regularity with its maximal Gröbner basis degree (i.e., the solving degree for the homogenized sequence). The definition of a generalized (cryptographic) semi-regular sequence is also given, and it derives a new cryptographic assumption to estimate the security of cryptosystems. From our experimental observation, we raise a conjecture and some questions related to this generalized semi-regularity. These definitions and our results provide a theoretical formulation of (somehow heuristic) discussions done so far in the cryptographic community.
@@ 1 Introduction
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. We denote by $\mathbb{A}_{K}^{n}$ (resp. $\mathbb{P}_{K}^{n}$) the $n$-dimensional affine (resp. projective) space over $K$. Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over $K$. For a given monomial ordering $\prec$ on (the set of monomials in) $R$, let $\mathrm{LM}(f)$ denote the leading monomial of $f\in R\smallsetminus\{0\}$ with respect to it. For a non-empty subset $F\subset R\smallsetminus\{0\}$, put $\mathrm{LM}(F):=\{\mathrm{LM}(f):f\in F\}$. A set $F$ (resp. a sequence $\bm{F}$) of polynomials in $R$ is said to be homogeneous if the elements of $F$ (resp. $\bm{F}$) are all homogeneous, and otherwise $F$ is said to be affine. We denote by $\langle F\rangle_{R}$ (or $\langle F\rangle$ simply) the ideal generated by a non-empty subset $F$ of $R$. For a polynomial $f$ in $R\smallsetminus\{0\}$, let $f^{\rm top}$ denote its maximal total degree part which we call the top part of $f$, and let $f^{h}$ denote its homogenization in $R^{\prime}=R[y]$ by an extra variable $y$, see Subsection A.2 below for details. For a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus\{0\})^{m}$, we also set $\bm{F}^{\rm top}:=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ and $\bm{F}^{h}:=(f_{1}^{h},\ldots,f_{m}^{h})$. For a finitely generated graded $R$-(or $R^{\prime}$-)module $M$, we also denote by ${\rm HF}_{M}$ and ${\rm HS}_{M}$ its Hilbert function and its Hilbert–Poincaré series, respectively.
@@
A Gröbner basis of an ideal $I$ in $R$ is defined as a special kind of generating set for $I$, and it gives a computational tool to determine many properties of $I$. A typical application of computing Gröbner bases is solving the multivariate polynomial (MP) problem: Given a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of $m$ polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus\{0\}$, find $(a_{1},\ldots,a_{n})\in K^{n}$ such that $f_{i}(a_{1},\ldots,a_{n})=0$ for all $i$ with $1\leq i\leq m$. A particular case where $f_{1},\ldots,f_{m}$ are all quadratic is called the MQ problem (cf. [44]), and its hardness is applied to constructing public-key cryptosystems that are expected to be quantum resistant. Therefore, analyzing the complexity of computing Gröbner bases is one of the most important problems both in theory and in practice.
@@
An algorithm for computing Gröbner bases was proposed first by Buchberger [7], and so far a number of its improvements such as the $F_{4}$ [19] and $F_{5}$ [20] algorithms have been proposed. In determining the complexity of computing Gröbner bases, as we will see in the first paragraph of Subsection 2.2 below, one of the most important cases is the case where the input system is zero-dimensional (see Terminology below for the meaning of zero-dimensional) and the monomial ordering is graded (i.e., degree-compatible). Therefore, we focus on that case in the rest of this paper. Namely, we suppose that the input sequence $\bm{F}=(f_{1},\ldots,f_{m})$ admits a finite number of zeros in $\mathbb{A}^{n}_{\overline{K}}$ (resp. $\mathbb{P}^{n-1}_{\overline{K}}$) if $\bm{F}$ is affine (resp. homogeneous), and we consider a monomial ordering $\prec$ on $R$ that compares monomials first by their total degrees, e.g., a degree reverse lexicographical (DRL) ordering. Then, the complexity of the Gröbner basis computation for $F=\{f_{1},\ldots,f_{m}\}$ is estimated as a function of the solving degree(s): To the authors’ best knowledge, there are three (in fact four) kinds of definitions of solving degree, and they will be rigorously described in Subsection 2.2 below. In the first definition, the solving degree is defined as the highest degree of the polynomials involved during the Gröbner basis computation. Since this solving degree depends on an algorithm $\mathcal{A}$ that one adopts, we denote it by $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$. On the other hand, in the second and the third definitions, which were originally provided in a series of Gorla et al.’s studies (cf. [9], [5], [25], [10], [24]), we can see that the solving degrees do not depend on an algorithm, but only on $F$ and $\prec$. The solving degree in the second (resp. third) definition is defined by using Macaulay matrices (resp. those with mutants), and it is denoted by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ (resp. $\mathrm{sd}_{\prec}^{\rm mut}(F)$) in this paper, where the subscripts “mac” and “mut” stand for Macaulay matrices and mutants respectively. Note that, when $F$ is homogeneous, these three solving degrees coincide with one another (for $\mathcal{A}$ with suitable setting) and we call them the solving degree simply; they are equal to the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec}(F)$ of $F$ with respect to $\prec$. In this case, we can apply a well-known bound [32, Theorem 2] by Lazard. In the following, we mainly treat with the case where $F$ is affine.
@@
  • ([9, Theorem 11]) When $K=\mathbb{F}_{q}$, the solving degree $\mathrm{sd}_{\prec}^{\rm mac}(F)$ for a DRL ordering $\prec$ can be upper-bounded by the Macaulay bound $d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell=\min\{n+1,\ell\}$, if $F$ contains the field equations $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
@@
  • Semaev-Tenti [41] (see also Tenti’s PhD thesis [42]) constructed a Buchberger-like algorithm $\mathcal{A}$ for the case $K=\mathbb{F}_{q}$ such that $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$ with $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$ for a DRL ordering $\prec$, assuming that $F=\{f_{1},\ldots,f_{m}\}\cup\{x_{i}^{q}-x\mid 1\leq i\leq n\}$ and $\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D$. Here $d_{\rm reg}(\langle F^{\rm top}\rangle)$ is the degree of regularity of $\langle F^{\rm top}\rangle$, i.e., the smallest non-negative integer $d$ with $R_{d}=\langle F^{\rm top}\rangle_{d}$, where $R_{d}$ denotes the homogeneous part (component) of degree $d$ and where we set $I_{d}=I\cap R_{d}$ for a homogeneous ideal $I$ of $R$.
  • Caminata-Gorla proved in [10, Theorem 3.1] that $\mathrm{sd}_{\prec}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\prec}}(F)\}$ with $F=\{f_{1},\ldots,f_{m}\}$ for any graded monomial ordering $\prec$, where $d_{F}$ denotes the last fall degree of $F$ defined in [10, Definition 1.5] (originally in [28], [27]). Recently, Salizzoni [40] also proved $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq D+1$, in the case where $\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D<\infty$.
In this paper, by a mathematically rigorous way following Gorla et al.’s works, we study the solving degrees and related Gröbner bases of affine semi-regular polynomial sequences, where a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of (not necessarily homogeneous) polynomials is said to be affine semi-regular (resp. affine cryptographic semi-regular) if $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is semi-regular (resp. cryptographic semi-regular), see Definitions 2.1.3, 2.1.10, and 2.1.13 for details. Note that homogeneous semi-regular sequences are conjectured by Pardue [37, Conjecture B] to be generic sequences of polynomials (see e.g., [37] for the definition of genericness), and affine (cryptographic) semi-regular sequences are often appearing in the construction of multivariate public key cryptosystems. As a sequel of the authors’ previous work [31], we investigate further results on the solving degrees and on behaviors of the computation of Gröbner bases.
@@
Theorem 1 (Theorem 3.1.1, Remark 3.1.2, and Corollary 3.1.4).
@@
    There exists $d_{0}$ such that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for all $d$ with $d\geq d_{0}$, namely the number of projective zeros for $F^{h}$ is finite at most.
@@
    ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\prod_{i=1}^{m}(1-z^{d_{i}})/(1-z)^{n+1}\pmod{z^{D}}$, so that $\bm{F}^{h}$ is $D$-regular, equivalently $\mathrm{syz}(F^{h})_{<D}=\mathrm{tsyz}(F^{h})_{<D}$. Here we denote by $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$ the module of syzygies of $F^{h}$ and that of trivial syzygies of $F^{h}$, respectively (see Appendix A.1 for the definition of $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$).
As for (3) of Theorem 1, similarly to the proof of Lazard’s bound [32, Theorem 2], it can be proved (see Proposition 2.3.6 below) that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\mathrm{max}\{D,D^{\prime}\}$, where $\prec^{h}$ is the homogenization of $\prec$ and where $D^{\prime}:=\min\{d_{0}\mid{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)\mbox{ for all $d\geq d_{0}$}\}$. As in [9, Theorem 11] recalled above, we can apply Lazard’s bound to obtaining $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{i}=\deg(f_{i})$ and $\ell=\min\{n+1,\ell\}$, assuming $d_{1}\geq\cdots\geq d_{m}$ in descending order. As an additional result in this paper, we also obtain the following upper-bound on the solving degree of $F^{h}$:
@@
    Let $S_{0}$ be the saturation exponent of $(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$, that is, the minimum non-negative integer $s$ such that $(\langle F^{h}\rangle:\langle y^{s}\rangle)=(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$. Then the solving degree of $F^{h}$ is upper-bounded by $D+{S_{0}}$.
Based on Theorem 1, we can explore the computations of reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ in Section 4 below, dividing the cases into the degree less than $D$ or not. More precisely, denoting by $G$, $G_{\rm hom}$, and $G_{\rm top}$ the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ respectively, where their monomial orderings are DRL $\prec$ or its homogenization $\prec^{h}$, we revisit [31, Section 5] and obtain more precise results:
Theorem 3 (Section 4; cf. [31, Section 5]).
@@
Remark 1.
In (4) of Theorem 3, the complexity of the algorithm $\mathcal{A}$ (with respect to the number of arithmetic operations on $K$) is
$O\left(m\binom{n+D}{D}^{\omega}+\binom{n+D}{D}^{2}\binom{n+D-1}{D-1}^{2}\binom{n+2D-2}{2D-2}\right),$
which is reduced to
$O\left(m\binom{n+D}{D}^{\omega}+\binom{n+D-1}{D-1}^{2}\binom{n+2D-2}{2D-2}\right)$
if we can avoid every 0-reduction completely, where $32\leq\omega<$ 3 is the exponent of matrix multiplication. This can be proved in a way similar to the proof of [42, Theorem 3.65] together with Theorem A.4.1 below: A proof will be given in Appendix A.5 below for the readers’ convenience.
@@
Furthermore, based on Lemma 2.2.2 below, for zero-dimensional homogeneous ideals (see Terminology below), we naturally extend the notion of degree of regularity: We shall define the generalized degree of regularity $\tilde{d}_{\rm reg}(I)$ of such a homogeneous ideal $I$, as the index of regularity (or called the Hilbert regularity) $i_{\rm reg}(I)$ of $I$. The generalized degree of regularity of $\langle F^{h}\rangle$ plays a very important role in analyzing the computation of Gröbner bases for such ideals, see Subsections 2.3 and 4.3. The following proposition summarizes several theoretical results, which are proved rigidly in this paper, on this generalized degree of regularity:
Proposition 1 (Lemma 2.3.5 and Proposition 2.3.6; see also Subsection A.3).
@@
    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial ordering $\prec$.
@@
Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [37, Section 4] for the original definition).
Corollary 1 (Corollary A.4.2 and Remark A.4.4; see also Subsection 4.3).
Under the same setting as in Proposition 1, when $d_{1},\ldots,d_{m}$ are fixed, a Gröbner basis of $\langle F\rangle_{R}$ can be computed in
$O\left(m\binom{n+\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)}{\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)}^{\omega}\right)$
if the complexity of substituting $y=1$ to $G_{\rm hom}$ is negligible.
Furthermore, if $m=n$, then $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)=d_{\rm reg}(\langle F^{\rm top}\rangle)-1$ with $d_{\rm reg}(\langle F^{\rm top}\rangle)=\sum_{j=1}^{n}d_{j}-n+1$. If $m>n$ and if $\bm{F}^{h}$ is generalized cryptographic semi-regular, i.e., $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$-regular, then we have
$\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\leq{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1,$
where []delimited-[][$\cdot$][ ⋅ ] means truncating a formal power series over $\mathbb{Z}$ after the last consecutive positive coefficient.
Finally, in Subsection 4.3, we give some observation on the behavior of Gröbner basis computation based on our experiments, from which we arrive at a conjecture (Conjecture 4.3.4 below) on the Hilbert–Poincaré series for affine polynomial sequences without constant terms. For this conjecture, we also generalize the notion of cryptographic semi-regular in Subsection 2.3.
@@
  • ${\rm HF}_{M}$: The Hilbert function of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HF}_{M}(d)=\mathrm{dim}_{K}M_{d}$ for each $d\in\mathbb{Z}$.
  • ${\rm HS}_{M}$: The Hilbert–Poincaré series of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}_{\geq 0}}M_{d}$, say ${\rm HS}_{M}(z)=\sum_{d=0}^{\infty}{\rm HF}_{M}(d)z^{d}\in\mathbb{Z}\llbracket z\rrbracket$.
@@
Terminology
Following [14, Chapter IX], we use the terminology ‘zero-dimensional’ for a polynomial ideal by the following rule:
  • If the ideal is not necessarily homogeneous, we say that it is zero-dimensional if the number of its affine zeros over $\overline{K}$ is finite at most. This is equivalent to that the affine Hilbert polynomial of the ideal is a constant.
  • If the ideal is homogeneous, unless otherwise noted, we say that it is zero-dimensional if the number of its projective zeros over $\overline{K}$ is finite at most. This is equivalent to that the Hilbert polynomial of the ideal is a constant.
As for the definition of affine Hilbert polynomials and Hilbert polynomials, we refer to [14, Chapter IX].